Phillips Pharmaceuticals Limited (PPL) shares are publicly traded on the Toronto Stock Exchange. The common shares currentlyHow can i calculate the cost of equity using Security Market Line and Dividend Discount Model (DDM), the cost of debt, the market value weights, and the WACC to use your answer from Security Market Line in your calculation for the cost of equity? Please do not use excel
Answer #1

COST OF EQUITY USING SECURITY MARKET LINE:

CAPM Equation:

Rs=Rf+Beta*(Rm-Rf)

Rs=Required Return

Rf=Risk Free Rate=4.25%

Beta=0.9

Rm=Expected Market Return=Expected Index Return=10%

Required Return =Rs=4.25%+0.9*(10-4.25)%=9.425%

FLOTATION COSTS=4.75%=0.0475

Cost of Equity =9.425/(1-0.0475)=9.9%

COST OF EQUITY USING DIVIDEND DISTRIBUTION MODEL:

Sustainable Growth Rate =(1-Payout Ratio)*Return on Equity

Payout ratio =60%=0.6

Return on Equity =10%

g=Sustainable Growth Rate=(1-0.6)*10%=4%=0.04

R=(D1/P0)+g

R=Required Return on Equity

D1=Next Years Dividend=D0*(1+g)

D0= Current Dividend=$1.50

D1=1.50*1.04=$1.56

P0=Current Market Price =$30.00

Required Return on Equity =(1.56/30)+0.04=0.092=9.2%

FLOTATION COSTS=4.75%=0.0475

Cost of Equity =9.2/(1-0.0475)=9.7%

COST OF DEBT:

You will need to use excel to get the cost of debt:It will not be able to calculate without using excel because of complicated equations involved.

Nper=Number of semi annual period left =10*2=20

Pmt = Semi annual Coupon Payment =(1000000*5.75%)/2=$28,750

Pv=Current Market Value =0.9654*1000000=$965,400

Fv=Amount to be paid on maturity after 10 years=$1000000

RATE=Semi annual yield to maturity=3.1% (Using RATE Function of excel with Nper=20, Pmt=28750,Pv=-965400,Fv=1000000)

Annual Yield to maturity =3.1*2=6.2%

The details of using Excel RATE function is attached.

Flotation Cost =3.67%=0.0367

Before tax cost of Debt =6.2/(1-0.0367)=6.4%

Tax Rate=40%=0.4

After Tax Cost of Debt=6.4*(1-0.4)=3.9%

MARKET VALUE WEIGHTS:

Equity =$30*100000 shares=$3,000,000

Debt =Pv=$965,400

Total Market Value=3000000+965400=$3,965,400

Market Value Weight of Equity =3000000/3965400=0.7565

Market Value Weight of Debt =965400/3965400=0.2435

WACC=Weight of Equity*Cost of Equity+Weight of debt*Cost of Debt

WACC=0.7565*9.9%+0.2435*3.9%=8.4%

Clipboard Font Alignment y Number D31 x fc =RATE(D27,028,-D29,030) E F ДА В Nper Pmt Pv FV RATE 32 Number of Semi annual coup